# DreamPRM: Domain-Reweighted Process Reward Model for Multimodal Reasoning

### v2

@@ 5.3 Benchmark evaluation of DreamPRM
o4-mini (4-16-25) [39] 80.6 81.6 82.5 84.2 85.2
Figure 5: Scaling ability and cross-model generalization. (a) Radar chart of five multimodal reasoning benchmarks shows that DreamPRM delivers monotonic accuracy gains as the number of selected chains-of-thought increases (@2, @4, @8) over the pass@1 baseline. (b) Best-of-N accuracy curves for InternVL-2.5-8B-MPO (blue) and GPT-4.1-mini (red) on MathVista confirm that the same DreamPRM-ranked CoTs generalize across models, consistently outperforming pass@1 performance (dashed lines) as $k$ grows.
@@
5.4 Scaling and generalization analysis of DreamPRM
DreamPRM scales reliably with more CoT candidates. As shown in the left panel of Fig. 5, the accuracy of DreamPRM consistently improves on all five benchmarks as the number of CoTs increases from $k{=}2$ to $k{=}8$, expanding the radar plot outward. Intuitively, a larger set of candidates increases the likelihood of including high-quality reasoning trajectories, but it also makes identifying the best ones more challenging. The consistent performance gains indicate that DreamPRM effectively verifies and ranks CoTs, demonstrating its robustness in selecting high-quality reasoning trajectories under more complex candidate pools.
DreamPRM transfers seamlessly to stronger base MLLMs. The right panel of Fig.5 shows the MathVista accuracy when applying DreamPRM to recent MLLMs, GPT-4.1-mini (2025-04-14) [40] and o4-mini (2025-04-16) [39]. For o4-mini model, the pass@1 score of 80.6% steadily increases to 85.2% at $k{=}8$, surpassing the previous state-of-the-art performance. This best-of-$N$ trend, previously observed with InternVL, also holds for GPT-4.1-mini and o4-mini, demonstrating the generalization ability of DreamPRM. Full results of these experiments are provided in Tab. 2.
5.5 Ablation study
In this section, we investigate the importance of three components in DreamPRM: (1) bi-level optimization, (2) aggregation function loss in upper-level, and (3) structural thinking prompt (detailed in Section 5.2). As shown in the rightmost panel of Fig. 4, the complete DreamPRM achieves the best results compared to three ablation baselines across all five benchmarks. Eliminating bi-level optimization causes large performance drop (e.g., -3.5% on MathVista and -3.4% on MMStar). Removing aggregation function loss leads to a consistent 1%-2% decline (e.g., 57.4% $\rightarrow$ 56.3% on WeMath). Excluding structural thinking also degrades performance (e.g., -1.8% on MathVision). These results indicate that all three components are critical for DreamPRM to achieve the best performance. More detailed results are shown in Tab. 3.
5.6 Analysis of learned domain weights
6 Conclusions
Appendix A Optimization algorithm
Appendix B Dataset Details
Task Dataset
Science AI2D [21], ScienceQA [32], M3CoT [5]
Chart ChartQA [36], DVQA [18], MapQA [2], FigureQA [19]
Commonsense IconQA [33], InfographicsVQA [37], CLEVR-Math [27]
To enhance the robustness of DreamPRM, we collect a diverse set of datasets in lower-level optimization, spanning multiple domains to ensure a comprehensive coverage of multimodal reasoning tasks, as reported in Tab. 4. The selected 15 multimodal datasets covers 4 major categories including science, chart, geometry and commonsense, with a wide range of task types (QA, OCR, spatial understanding). Additionally, we observe that for some questions, given the current structural thinking prompts, MLLMs consistently produce either correct or incorrect answers. Continuing to sample such questions is a waste of computational resources. Inspired by the dynamic sampling strategy in DAPO [69], we propose a similar dynamic sampling technique for Monte Carlo estimation that focuses on prompts with varied outcomes to improve efficiency. After processing and sampling, the training datasets in lower-level $\mathcal{D}_{tr}$ have around 15k examples (1k per each of the 15 domains), while the meta dataset in the upper-level $\mathcal{D}_{meta}$ has around 1k validation examples from the MMMU [70] dataset.
Appendix C Structural Thinking Prompt
Figure 7: Zero-shot prompting for structural thinking.
@@ Appendix D Additional Experimental Results
Loss curves and domain weights. The loss curves and domain weights during the fine-tuning of DreamPRM are illustrated in Fig. 8. It can be observed that the learnt distribution emphasizes informative mathematical figure domains while attenuating less relevant sources. Additionally, domain weights start at 1.0 and quickly diverge, stabilizing after roughly half the training, and the inner and outer losses decrease steadily and plateau, indicating stable convergence of the bi-level training procedure.

### v3

@@ Appendix D Additional Experimental Results
Leaderboard performance details. Table 5 presents a comprehensive comparison of different PRM variants built upon the same o4-mini backbone. DreamPRM consistently outperforms all baselines, elevating the base o4-mini performance from 80.6These steady improvements demonstrate the effectiveness of DreamPRM in enhancing reasoning accuracy through process-level supervision and promoting more reliable consensus across multiple chains of thought.
Best-of-N results. Tab. 3 reports the accuracy of two state-of-the-art models on MathVista dataset using DreamPRM with varying numbers $k$ of CoTs. The results indicate that the performance scales well with the number of CoTs.
Table 5: Comparison of different PRM variants on the o4-mini model (evaluated on eight CoTs).
Method Accuracy
o4-mini 80.6
   + Self-consistency 81.7
   + ORM 80.8
   + Vanilla-PRM 84.2
   + DreamPRM 85.2
Ablation studies. The exact results of ablation experiments in the main paper are included in Tab. 4, which emphasizes the importance of all the components in DreamPRM.
Loss curves and domain weights. The loss curves and domain weights during the fine-tuning of DreamPRM are illustrated in Fig. 9. It can be observed that the learnt distribution emphasizes informative mathematical figure domains while attenuating less relevant sources. Additionally, domain weights start at 1.0 and quickly diverge, stabilizing after roughly half the training, and the inner and outer losses decrease steadily and plateau, indicating stable convergence of the bi‑level training procedure.